### v1

@@ Abstract
Such a scheme has surprising benefits both in theory and practice. We leverage the approach to improve the representation of any move structure with $r$ runs over a domain $n$ to $O(r\log r+r\log\frac{n}{r})$-bits of space. The worst case query time is also improved to $O(\log\frac{n}{r})$ without balancing. An $O(r)$-time and $O(r)$-space construction lets us apply the method to run-length encoded $\mathrm{BWT}$ ($\mathrm{RLBWT}$) permutations such as $\mathrm{LF}$ and $\phi$ to obtain optimal-time algorithms for $\mathrm{BWT}$ inversion and suffix array ($\mathrm{SA}$) enumeration in $O(r)$ additional working space. Finally, we provide the RunPerm library, providing flexible plug and play move structure support, and use it to evaluate our splitting approach. Experiments find length capping results in faster move structures, but also a space reduction: at least $\sim 40\%$ for $\mathrm{LF}$ across large repetitive genomic collections.
@@ 1 Introduction
Given a permutation $\pi$ over domain $n$ with $r$ contiguously permuted runs (or intervals), we prove that by splitting intervals of length greater than a constant factor of the average length $\frac{n}{r}$, a length capped move structure is obtained that achieves constant-time per move query averaged over the $n$ distinct permutation steps in $O(r)$-space. Further, this can be done in $O(r)$-time in comparison to the $O(r\log r)$-time required for balancing [bertram2024move]. Surprisingly, this also obtains an $O(r\log r+r\log\frac{n}{r})$-bit representation for any move structure, including when used alongside balancing, which improves upon the existing $O(r\log n)$-bit result. A consequence of length capping also bounds the worst case query to $O(\log\frac{n}{r})$ as an alternative to the existing $O(\log r)$-time of an unbalanced representation.
@@ 3.2 Length Capping
3.3 Other Theoretical Improvements
A move structure maintaining $O(1)$-time move queries can be represented by $S$ as a sparse bitvector using $O(r+r\log\frac{n}{r})$-bits, $S_{\pi}$ in $r\log n$ bits, and $S_{rank(\pi)}$ in $r\log r$ bits. Overall, this representation results in $O(r\log n)$-bits. Switching $S_{\pi}$ for $S_{\Delta}$ may be smaller in practice, but is still worst case $O(r\log n)$-bits. However, after length capping, since $S_{\Delta}$ is the offset within an interval, it is upper bounded by the maximum length $L\in O(\frac{n}{r})$. Hence, we can represent a move structure in overall $O(r\log r+r\log\frac{n}{r})$-bits. Further, if we perform length capping first and then perform balancing, we get theoretical space improvements combined with worst case guarantees. Also note that we can replace $S$ with $S_{\ell}$ when working with relative positions and achieve the same bounds, which may be useful in practice since array access is more cache efficient than sparse bitvector access.
@@ 3.3 Other Theoretical Improvements
Given a permutation $\pi$ over [0,n1][0,n-1], let $S$ be the set of $r$ positions such that either $i=0$ or $\pi(i-1)\neq\pi(i)-1$. We can construct in $O(r\log r)$-time and $O(r)$-space a length capped and balanced move structure, which, given position $i$ and the rank of its predecessor in $S$, computes a move query returning $\pi(i)$ and the rank of $\pi(i)$’s predecessor in $S$ in $O(1)$-time and $O(r\log r+r\log\frac{n}{r})$-bits of space.
@@
We find $O(r)$ intervals by length capping and then perform balancing, which maintains $O(r)$ intervals after completion. Combined with Theorem 2.1 and the representation of the move structure using $S_{\Delta}$ in $O(r\log\frac{n}{r})$-bits we achieve the desired result.
@@ 5.1 RunPerm Library
Beyond a proof of concept for length capping, RunPerm is designed as a header only library that can be inserted into existing software. As such, it supports various representations of a move structure. For absolute positions, the components $S$, $S_{\Delta}$ and $S_{rank(\pi)}$ are stored in a bitpacked matrix using $O(r\log n)$-bits. Alternatively, relative positions are supported by storing the components $S_{\ell}$, $S_{\Delta}$, and $S_{rank(\pi)}$ in $O(r\log r+r\log\frac{n}{r})$-bits as described in Section 3.3. Move queries support using either a traditional linear scan over fast forwards or exponential search as in Corollary 3.6.
@@ 6 Conclusion
In this work, we showed an additional splitting regime which achieves average time guarantees with respect to move queries and amortized time guarantees for single cycle permutations while obtaining a $O(r)$-time construction superior to that of balancing. This can be leveraged in many $\mathrm{RLBWT}$ permutations such as $\mathrm{LF}$ and $\phi$ to obtain new optimal-time results in $\mathrm{BWT}$ inversion and $\mathrm{SA}$ enumeration. Further, length capping achieves an improved $O(r\log r+r\log\frac{n}{r})$-bit move structure representation and new worst case $O(\log\frac{n}{r})$-time when not explicitly balancing.
@@
In essence, length capping can be seen as “fixing” a distribution whose values sum to at most $n$ even if a single value can be $O(n)$ itself; this is true of both $S_{\ell}$ and $S_{\Delta}$. It is likely that such an approach could be adopted in other such scenarios. Further, the unique properties of $\mathrm{RLBWT}$ based permutations allow efficient unbalanced move structure construction. It would seem that most runny permutations would have similar exploitable structures, since understanding their underlying properties appears necessary to derive their function in the first place.

### v2

@@ Abstract
Such a scheme has surprising benefits both in theory and practice. For a move structure with $r$ runs over a domain $n$, we replace all $O(r\log n)$-bit components to reduce the overall representation by $O(r\log r)$-bits. The worst case query time is also improved to $O(\log\frac{n}{r})$ without balancing. An $O(r)$-time and $O(r)$-space construction lets us apply the method to run-length encoded $\mathrm{BWT}$ ($\mathrm{RLBWT}$) permutations such as $\mathrm{LF}$ and $\phi$ to obtain optimal-time algorithms for $\mathrm{BWT}$ inversion and suffix array ($\mathrm{SA}$) enumeration in $O(r)$ additional working space. Finally, we provide the RunPerm library, providing flexible plug and play move structure support, and use it to evaluate our splitting approach. Experiments find length capping results in faster move structures, but also a space reduction: at least $\sim 40\%$ for $\mathrm{LF}$ across large repetitive genomic collections.
@@ 1 Introduction
Given a permutation $\pi$ over domain $n$ with $r$ contiguously permuted runs (or intervals), we prove that by splitting intervals of length greater than a constant factor of the average length $\frac{n}{r}$, a length capped move structure is obtained that achieves constant-time per move query averaged over the $n$ distinct permutation steps in $O(r)$-space. Further, this can be done in $O(r)$-time in comparison to the $O(r\log r)$-time required for balancing [bertram2024move]. This also allows us to replace all $O(r\log n)$ bit components of an arbitrary move structure with $O(r\log\frac{n}{r})$ bit representations. Overall, this saves $O(r\log r)$ bits of total space. A consequence of length capping also bounds the worst case query to $O(\log\frac{n}{r})$ as an alternative to the existing $O(\log r)$-time of an unbalanced representation.
@@ 3.2 Length Capping
3.3 Other Results
A move structure maintaining $O(1)$-time move queries can be represented by $S$ as a sparse bitvector using $r\log\frac{n}{r}+O(r)$-bits, $S_{\pi}$ in $r\log n$ bits, and $S_{rank(\pi)}$ in $r\log r$ bits. Overall, this representation results in $O(r\log n)$-bits. Although Brown et al. [brown2024faster] showed that move structures for $\mathrm{LF}$ and $\mathrm{FL}$ can be represented in $O(r\log\frac{n}{r})$-bits, we do not know how to remove the $r\log n$ bit components for general permutations, even when switching $S_{\pi}$ for $S_{\Delta}$.
However, after length capping, since any element in $S_{\Delta}$ is the offset within an interval, it is upper bounded by the maximum length $L\in O(\frac{n}{r})$. Thus, with length splitting, we can support both absolute and relative positions by representing $S_{\Delta}$ in $O(r\log\frac{n}{r})$-bits. If using relative positions, since any element in $S_{\ell}$ is also bounded by $L$, we can represent it in the same $O(r\log\frac{n}{r})$-bits as representing $S$ using a sparse bitvector. This may be useful in practice since array access is cache friendly compared to sparse bitvector select queries.
Overall, this results in $O(r\log r+r\log\frac{n}{r})=O(r\log n)$-bits, so the asymptotic bound is unchanged. However, no single component requires $O(r\log n)$-bits. Let $\hat{r}=\Theta(r)$ to suppress constant factors of $r$ introduced by balancing and/or length capping. Then the total space in bits by applying length capping changes from
[…]
to
[…]
effectively saving $O(r\log r)$-bits of space. This result can also be applied by length capping alongside balancing.
@@ 3.3 Other Results
[…]
@@
[…]
@@ 5.1 RunPerm Library
Beyond a proof of concept for length capping, RunPerm is designed as a header only library that can be inserted into existing software. As such, it supports various representations of a move structure. For absolute positions, the components $S$, $S_{\Delta}$ and $S_{rank(\pi)}$ are stored in a bitpacked matrix using roughly $r\log r+2r\log n$ bits (we do not use sparse bitvectors, optimizing instead for speed). Alternatively, relative positions are supported by storing the components $S_{\ell}$, $S_{\Delta}$, and $S_{rank(\pi)}$ in roughly $r\log r+2r\log\frac{n}{r}$ bits as described in Section 3.3. Move queries support using either a traditional linear scan over fast forwards or exponential search as in Corollary 3.6.
@@ 6 Conclusion
In this work, we showed an additional splitting regime which achieves average time guarantees with respect to move queries and amortized time guarantees for single cycle permutations while obtaining a $O(r)$-time construction superior to that of balancing. This can be leveraged in many $\mathrm{RLBWT}$ permutations such as $\mathrm{LF}$ and $\phi$ to obtain new optimal-time results in $\mathrm{BWT}$ inversion and $\mathrm{SA}$ enumeration. Further, length capping achieves an $O(r\log r)$-bit space reduction of arbitrary move structures by replacing all $O(r\log n)$-bit components. Finally, we provide a new worst case move query bound for unbalanced move structures of $O(\log\frac{n}{r})$-time.
@@
In essence, length capping can be seen as “fixing” a distribution whose values sum to at most $n$ even if a single value can be $O(n)$ itself; this is true of both $S_{\ell}$ and $S_{\Delta}$. It is likely that such an approach could be adopted in other such scenarios. This result could also be combined with Elias-Fano encoding, since lengths store the lower $\log\frac{n}{r}$ bits of elements in $S$ in a cache efficient manner which could be augmented with a bitvector for the higher bits. Further, the unique properties of $\mathrm{RLBWT}$ based permutations allow efficient unbalanced move structure construction. It would seem that most runny permutations would have similar exploitable structures, since understanding their underlying properties appears necessary to derive their function in the first place.